# Photodissociation spectra of single trapped CaOH molecular ions

### v1

@@ Abstract
Molecular ions that are generated by chemical reactions with trapped atomic ions can serve as an accessible and successful testbed for developing molecular quantum technologies. On the other hand, they are also a hindrance to scaling up quantum computers based on atomic ions as unavoidable reactions with background gas destroy the information carriers. Here, we investigate the single-photon and two-photon dissociation processes of single CaOH molecular ions co-trapped in Ca ion crystals using a femtosecond laser system. We report the photodissociation cross section spectra of CaOH for single-photon processes at nm and for two-photon processes at nm. This result can serve as a basis for dissociation-based spectroscopy for studying the internal structure of CaOH. The result also gives a prescription for recycling Ca ions in large-scale trapped Ca quantum experiments from undesired CaOH ions formed in the presence of background water vapor.
@@ I Introduction
For trapped ion experiments based on Ca, which is one of the most well-studied and widely implemented ion species for quantum information processingSchindler et al. (2013), molecular ions generated in chemical reactions with background gas are predominantly CaH, CaO, and CaOHHansen et al. (2012). While extensive theoretical and experimental research has been conducted on the two diatomic speciesAbe et al. (2012); Rugango et al. (2016); Chou et al. (2017); VanGundy, Bartlett, and Heaven (2018); Chou et al. (2020); Qi, Reed, and Brown (2023), there exists little spectroscopic study concerning the triatomic CaOH molecular ion.
As illustrated in Fig. 1.(a), CaOH can form in a trapped Ca ion experiment in the presence of water vapor in the vacuum chamber by the reactionOkada et al. (2003)
While CaH and CaO can be photodissociated with the 397 nm and 375 nm lasers used in trapped Ca experimentsWu et al. (2021), the photodissociation of CaOH requires a shorter wavelength for exciting the molecule to its unbound first electronic excited state. This photodissociation channel was studied by Okada et al. in 2006 using a combination of optical filters and a broadband UV lamp, where they estimated the dissociation threshold to be Okada et al. (2006). However, it was later reported by Hansen et al. that Ca can also be recovered from CaOH by applying 272 nm lightHansen et al. (2012). In this paper, we report on a measurement of the photodissociation spectrum of CaOH, where the wavelength-dependence of the photodissociation cross section was measured by excitation with a femtosecond optical parametric amplifier (OPA) system. With the same ultrafast laser system, we are also able to investigate two-photon dissociation of CaOH and the measured two-photon dissociation spectrum is reported. The result implies the possibility of recycling Ca from CaOH with ablation lasers at 515 nm, which are commonly used in trapped Ca quantum experiments.
@@ II Experimental setup
As illustrated in Fig. 2, our experiment is based on a linear Paul trapPogorelov et al. (2021) in ultrahigh vacuum (UHV). A gate valve separates the chamber into two parts: the main experiment chamber and the molecular gas chamber. The main experiment chamber containing the ion trap is connected to a non-evaporable getter (NEG) pump and a NEG-ion combination pump to maintain a base pressure mbar. Ca ions in the ion trap originate from an integrated Ca target mounted near the trap. By ablating the target with 532 nm laser pulses, neutral Ca atoms are introduced to the ion trap where they are photoionized in a two-step process using 375 nm and 422 nm lasersPogorelov et al. (2021). The generated Ca ions are then captured by the ion trap and Doppler-cooled by driving their 4S 4P cycling transition at 397 nm, as shown in the inset of Fig. 2. Additional lasers at 866 nm and 854 nm are applied to repump the ions that decay to the 3D and 3D states, respectively, back to the cyclic cooling manifold.
To the right of the main chamber, the molecular gas chamber is used to accumulate water vapor for the generation of CaOH molecular ions. When the gate valve is closed, this section is not pumped by the vacuum pumps attached to the main chamber and the water vapor pressure inside increases over time due to outgassing from the chamber wall. To increase the partial pressure of water further, water vapor that is outgassing from an unbaked stainless steel hose is introduced to the molecular gas chamber via a leak valve.
@@ III Molecule Generation and identification
To generate CaOH for the photodissociation measurement, water vapor is introduced to the main chamber by opening the gate valve. In our setup, the partial pressure of water vapor in the system cannot be directly measured. Instead, the increase in water vapor pressure in the main chamber is inferred from the change in total pressure estimated from the current of the ion pump. While waiting for the generation of CaOH, the gate valve is opened every 5 minutes for 1 second, which results in the total pressure estimated by the ion pump increasing from mbar to mbar.
As investigated by Okada et al. in 2003, the chemical reaction rate between HO and the trapped Ca ions increases when the ions are pumped into the 3D state by switching off the 866 nm repumping light during laser coolingOkada et al. (2003). Therefore, the 866 nm laser is periodically switched off for 1 second every 2 seconds in the molecule generation phase. The generation rate of one CaOH molecular ion is typically 10 to 15 minutes for the ion crystals of Ca ions investigated in this experiment.
The above molecule generation process was found to produce predominantly CaOH by mass spectrometry using the method presented by Drewsen et al. Drewsen et al. (2004) (for details, refer to Appendix D). In this method, the unknown molecular mass can be inferred from the frequency of the axial center-of-mass (COM) motion of a 2-ion crystal of one Ca ion and one molecular ionKielpinski et al. (2000); Drewsen et al. (2004) as
where is the ratio between the atomic and molecular mass and is the COM motional frequency of a Ca ion crystal. These frequencies were measured by applying an oscillating voltage to one of the endcap electrodes. When the frequency of this “tickling” signal is resonant with that of the COM motion of the ion crystal, it excites the corresponding oscillation of the crystal. This increases the motional temperature of the ion crystal and thus reduces the scattered Doppler cooling light. The COM motional frequency is thus determined from the ion crystal’s fluorescence as a function of the tickling frequency. The masses of the generated molecular ions are measured with a precision of m amu and the results are shown in Fig. 3. All 65 measurements have masses consistent with CaOH, from which we conclude that the molecular ions generated in our procedure are predominantly this species.
@@ IV Photodissociation
After verifying the generation of a single CaOH, the measurement of the photodissociation spectrum of CaOH is performed by applying femtosecond pulse trains produced by an OPA. The spatial beam profile and position of the OPA light are measured by a beam profiler mounted on a rail. To assist the alignment of the OPA light, an additional 397 nm Doppler cooling beam is applied co-propagating with the OPA light, which is aligned to the center of the ion trap by maximizing the 397 nm light scattered by the ion crystal. The OPA light is overlapped with this axial Doppler cooling beam at the position of the ions. We estimate the alignment precision of this procedure to be 10 µm, limited by the resolution of the beam profiler. The radius of the OPA light at the ion crystal was also estimated by the beam profiler to be 150(10) µm.
We assume that the dissociation of a CaOH molecule during illumination is memoryless due to the lack of intermediate electronic states. Thus, the dissociation probability after illumination time can be described by the cumulative distribution function (CDF) of an exponential distribution,
@@
The experimental procedure is initialized by loading a one-dimensional ion chain with Ca ions into the ion trap for the generation of a single CaOH (see Fig. 10). While Ca ions can be detected on an electron-multipling CCD (EMCCD) camera, the generated CaOH ions do not scatter light and are recognized as dark defects in the ion chain. The number of dark ions is identified by comparing the ion crystal image with pre-calculated ion crystal configuration templates with different numbers of dark and bright ions (see Appendix C). After the appearance of a single molecular ion, the OPA light is applied to illuminate the molecular ion for up to 300 s. If the OPA light successfully recovers the original Ca from the CaOH ion, as shown in Fig. 4, the illumination time for achieving photodissociation will be recorded. Otherwise, if the photodissociation is not successful or the dark ion is lost during the process, the dissociation trial is marked as unsuccessful within the accumulated illumination time. In either case, utilizing Bayesian analysis, the -th trial provides information to update the prior distribution of the mean dissociation time by multiplying it with a likelihood function
@@
reveals an estimation of . To ensure convergence of the posterior distribution, approximately 20 trials were performed at each wavelength.
@@ V Two-photon dissociation
In addition to the measurement of single-photon dissociation, we also investigated the photodissociation of CaOH at wavelengths around 500 nm. To achieve a dissociation rate that is comparable to the single-photon measurements, the repetition rate of the OPA is increased to 10 kHz and the light intensity is increased by reducing the beam radius to 70(10) µm and increasing the average power to a few milliwatts. At nm, the dependence of the dissociation rate on the average intensity of the OPA light on the ions is investigated, as illustrated in Fig. 6.(a). Fitting the data to a power law function gives an exponent of , which is consistent with two-photon processes.
Thus, the photodissociation of CaOH at wavelengths around 500 nm is assumed to be dominated by a two-photon dissociation channel via its unbound first electronic excited state. The two-photon dissociation cross section is related to the dissociation rate by (see Appendix A.2)
@@
VI Discussion
@@
In summary, we measured the photodissociation spectrum of single trapped CaOH molecular ions sympathetically cooled in Ca ion crystals at nm and nm. The photodissociation processes at wavelengths in these two ranges are verified to originate predominantly from a single-photon and a two-photon dissociation channel, respectively. While the one-photon dissociation cross section started to drop off around nm, a similar decrease was observed in the two-photon dissociation spectrum around nm. This implies that both processes involve the excitation of the molecular ion to its unbound first electronic excited state which leads to its dissociation. Compared to the previous workOkada et al. (2006) on single-photon dissociation of CaOH by Okada et al. in 2006, we report a more precise measurement of the dissociation cross sections at various wavelengths which is consistent with the previous estimation of at nm. The measured single-photon dissociation spectrum is also consistent with the photodissociation of CaOH at 272 nm observed by Hansen et al.Hansen et al. (2012).
@@
As for the two-photon dissociation of CaOH around nm, the dissociation process is found to be significantly less efficient compared to the single-photon dissociation process, as expected since higher peak intensity is typically required for higher-order processes. Yet femtosecond lasers at such wavelengths have proven to be alternatives to continuous wave lasers with nm for photodissociation of the molecular ion. The question arises as to whether nanosecond lasers that are used for ablation loading in trapped atomic ion experiments at nmPogorelov et al. (2021) are able to dissociate CaOH. Achieving a mean photodissociation time of 1 minute for a single CaOH ion is estimated to require an average power of W, considering a laser with kHz, ns, and the light focused onto the ions with a beam waist of µm.
@@
The reported measurement is implemented on single CaOH molecular ions in small trapped ion crystals, which enables the relatively precise evaluation of the photodissociation cross section due to the repeatable positioning of the molecule. The scale of the system is also crucial for investigating a two-photon dissociation process that requires tightly focused light to achieve sufficient intensity. The reported photodissociation spectrum can be a basis for studying the rovibrational structure of CaOH via dissociation-based spectroscopy. Designed to investigate the interaction between femtosecond laser pulses and a linear ion crystal with a single CaOH, this experiment represents a precursor for performing quantum logic spectroscopy with polyatomic molecular ions Schindler (2019).

### v2

# Photodissociation spectra of single trapped CaOH+ molecular ions

@@ Abstract
Molecular ions that are generated by chemical reactions with trapped atomic ions can serve as an accessible testbed for developing molecular quantum technologies. On the other hand, they are also a hindrance to scaling up quantum computers based on atomic ions as unavoidable reactions with background gas destroy the information carriers. Here, we investigate the single- and two-photon dissociation processes of single CaOH+ molecular ions co-trapped in Ca+ ion crystals using a femtosecond laser system. We report the photodissociation cross section spectra of CaOH+ for single-photon processes at nm and for two-photon processes at nm. Measurements are interpreted with quantum-chemical calculations, which predict the photodissociation threshold for at 265 nm. This result can serve as a basis for dissociation-based spectroscopy for studying the internal structure of CaOH+. The result also gives a prescription for recycling Ca+ ions in large-scale trapped Ca+ quantum experiments from undesired CaOH+ ions formed in the presence of background water vapor.
@@ I Introduction
For trapped ion experiments based on 40Ca+, which is one of the most well-studied and widely implemented ion species for quantum information processingSchindler et al. (2013), molecular ions generated in chemical reactions with background gas are predominantly CaH+, CaO+, and CaOH+Hansen et al. (2012). While extensive theoretical and experimental research has been conducted on the two diatomic speciesAbe et al. (2012); Rugango et al. (2016); Chou et al. (2017); VanGundy, Bartlett, and Heaven (2018); Chou et al. (2020); Qi, Reed, and Brown (2023), there are few spectroscopic studies concerning the triatomic CaOH+ molecular ion.
As illustrated in Fig. 1.(a), CaOH+ can form in a trapped Ca+ ion experiment in the presence of water vapor in the vacuum chamber by the reactionOkada et al. (2003)
While CaH+ and CaO+ can be photodissociated with the 397 nm and 375 nm lasers used in trapped Ca+ experimentsWu et al. (2021), the photodissociation of CaOH+ requires a shorter wavelength for exciting the molecule to its unbound first electronic excited state. This photodissociation channel was studied by Okada et al. in 2006 using a combination of optical filters and a broadband UV lamp, where they estimated the dissociation threshold to be Okada et al. (2006). However, it was later reported by Hansen et al. that Ca+ can also be recovered from CaOH+ by applying 272 nm lightHansen et al. (2012). In this paper, we report on a measurement of the photodissociation spectrum of CaOH+, where the wavelength-dependence of the photodissociation cross section was measured by excitation with a femtosecond optical parametric amplifier (OPA) system. With the same ultrafast laser system, we are also able to investigate two-photon dissociation of CaOH+ and the measured two-photon dissociation spectrum is reported. The result implies the possibility of recycling Ca+ from CaOH+ with ablation lasers at 515 nm, which are commonly used in trapped Ca+ quantum experiments. We employ quantum-chemical calculations to reproduce the measured photodissociation energy and predict the spectrum of excited electronic states of CaOH+.
@@ II Experimental setup
As illustrated in Fig. 2, our experiment is based on a linear Paul trapPogorelov et al. (2021) in ultrahigh vacuum (UHV). A gate valve separates the chamber into two parts: the main experiment chamber and the molecular gas chamber. The main experiment chamber containing the ion trap is connected to a non-evaporable getter (NEG) pump and a NEG-ion combination pump to maintain a base pressure mbar. Ca+ ions in the ion trap originate from an integrated 40Ca target mounted near the trap. By ablating the target with 532 nm laser pulses, neutral 40Ca atoms are introduced to the ion trap where they are photoionized in a two-step process using 375 nm and 422 nm lasersPogorelov et al. (2021). The generated Ca+ ions are then captured by the ion trap and Doppler-cooled by driving their 42S1/2 42P1/2 cycling transition at 397 nm, as shown in the inset of Fig. 2. Additional lasers at 866 nm and 854 nm are applied to repump the ions that decay to the 32D3/2 and 32D5/2 states, respectively, back to the cyclic cooling manifold.
To the right of the main chamber, the molecular gas chamber is used to accumulate water vapor for the generation of CaOH+ molecular ions. When the gate valve is closed, this section is not pumped by the vacuum pumps attached to the main chamber and the water vapor pressure inside increases over time due to outgassing from the chamber wall. To increase the partial pressure of water further, water vapor that is outgassing from an unbaked stainless steel hose is introduced to the molecular gas chamber via a leak valve.
@@ III Molecule Generation and identification
To generate CaOH+ for the photodissociation measurement, water vapor is introduced to the main chamber by opening the gate valve. In our setup, the partial pressure of water vapor in the system cannot be directly measured. Instead, the increase in water vapor pressure in the main chamber is inferred from the change in total pressure estimated from the current of the ion pump. While waiting for the generation of CaOH+, the gate valve is opened every 5 minutes for 1 second, which results in the total pressure estimated by the ion pump increasing from mbar to mbar.
As investigated by Okada et al. in 2003, the chemical reaction rate between H2O and the trapped Ca+ ions increases when the ions are pumped into the 32D3/2 state by switching off the 866 nm repumping light during laser coolingOkada et al. (2003). Therefore, the 866 nm laser is periodically switched off for 1 second every 2 seconds in the molecule generation phase. The generation rate of one CaOH+ molecular ion is typically 10 to 15 minutes for the ion crystals of Ca+ ions investigated in this experiment.
The above molecule generation process was found to produce predominantly CaOH+ by mass spectrometry using the method presented by Drewsen et al. Drewsen et al. (2004) (for details, refer to Appendix D). In this method, the unknown molecular mass can be inferred from the frequency of the axial center-of-mass (COM) motion of a 2-ion crystal of one Ca+ ion and one molecular ionKielpinski et al. (2000); Drewsen et al. (2004) as
where is the ratio between the atomic and molecular mass and is the COM motional frequency of a Ca+ ion crystal. These frequencies were measured by applying an oscillating voltage to one of the endcap electrodes. When the frequency of this “tickling” signal is resonant with that of the COM motion of the ion crystal, it excites the corresponding oscillation of the crystal. This increases the motional temperature of the ion crystal and thus reduces the scattered Doppler cooling light. The COM motional frequency is thus determined from the ion crystal’s fluorescence as a function of the tickling frequency. The masses of the generated molecular ions are measured with a precision of m amu and the results are shown in Fig. 3. All 65 measurements have masses consistent with CaOH+, from which we conclude that the molecular ions generated in our procedure are predominantly this species.
@@ IV Photodissociation
After verifying the generation of a single CaOH+, the measurement of the photodissociation spectrum of CaOH+ is performed by applying femtosecond pulse trains produced by an OPA. The spatial beam profile and position of the OPA light are measured by a beam profiler mounted on a rail. To assist the alignment of the OPA light, an additional 397 nm Doppler cooling beam is applied co-propagating with the OPA light, which is aligned to the center of the ion trap by maximizing the 397 nm light scattered by the ion crystal. The OPA light is overlapped with this axial Doppler cooling beam at the position of the ions. We estimate the alignment precision of this procedure to be 10 µm, limited by the resolution of the beam profiler. The radius of the OPA light at the ion crystal was also estimated by the beam profiler to be 150(10) µm.
We assume that the dissociation of a CaOH+ molecule during illumination is memoryless due to the lack of intermediate electronic states. Thus, the dissociation probability after illumination time can be described by the cumulative distribution function (CDF) of an exponential distribution,
@@
The experimental procedure is initialized by loading a one-dimensional ion chain with Ca+ ions into the ion trap for the generation of a single CaOH+ (see Fig. 11). While Ca+ ions can be detected on an electron-multipling CCD (EMCCD) camera, the generated CaOH+ ions do not scatter light and are recognized as dark defects in the ion chain. The number of dark ions is identified by comparing the ion crystal image with pre-calculated ion crystal configuration templates with different numbers of dark and bright ions (see Appendix C). After the appearance of a single molecular ion, the OPA light is applied to illuminate the molecular ion for up to 300 s. If the OPA light successfully recovers the original Ca+ from the CaOH+ ion, as shown in Fig. 4, the illumination time for achieving photodissociation will be recorded. Otherwise, if the photodissociation is not successful or the dark ion is lost during the process, the dissociation trial is marked as unsuccessful within the accumulated illumination time. In either case, utilizing Bayesian analysis, the -th trial provides information to update the prior distribution of the mean dissociation time by multiplying it with a likelihood function
@@
is normalized and reveals an estimation of . To ensure convergence of the posterior distribution, approximately 20 trials were performed at each wavelength.
@@ V Two-photon dissociation
In addition to the measurement of single-photon dissociation, we also investigated the photodissociation of CaOH+ at wavelengths around 500 nm. To achieve a dissociation rate that is comparable to the single-photon measurements, the repetition rate of the OPA is increased to 10 kHz and the light intensity is increased by reducing the beam radius to 70(10) µm and increasing the average power to a few milliwatts. At nm, the dependence of the dissociation rate on the average intensity of the OPA light on the ions is investigated, as illustrated in Fig. 6.(a). Fitting the data to a power law function gives an exponent of , which is consistent with two-photon processes.
Thus, the photodissociation of CaOH+ at wavelengths around 500 nm is assumed to be dominated by a two-photon dissociation channel via its unbound first electronic excited state. The two-photon dissociation cross section is related to the dissociation rate by (see Appendix A.2)
@@
VI Quantum-chemical modelling
To interpret experimental measurements and predict theoretical spectroscopic parameters of CaOH+, we use state-of-the-art quantum-chemical methods. We employ wave-function approaches: the coupled cluster method restricted to single, double, and noniterative triple excitations, CCSD(T) Hampel, Peterson, and Werner (1992), to describe the ground electronic state and internally contracted multireference configuration interaction method restricted to single and double excitations, MRCISD Werner and Knowles (1988), to describe excited electronic states. Electronic orbitals are constructed using the augmented correlation-consistent polarized weighted core-valence quintuple- quality basis set (aug-cc-pwCV5Z-PP Hill and Peterson (2017) for Ca, aug-cc-pwCV5Z Peterson and Dunning (2002) for O, and aug-cc-pV5Z Dunning (1989) for H). We include the scalar relativistic effect in Ca using the small-core relativistic energy-consistent pseudopotential (ECP10MDF Lim, Stoll, and Schwerdtfeger (2006)) to replace the inner-shell electrons. Electronic structure calculations are performed with the Molpro Werner et al. (2012a, b) package of ab initio programs. Uncertainties of present energy calculations should not be larger than 1-2%Gronowski, Koza, and Tomza (2020).
We optimize the ground-state geometry of linear CaOH+ and find the equilibrium internuclear distances bohr and bohr. Such an equilibrium geometry gives the rotational constant cm-1. Present theoretical harmonic frequencies for normal modes of CaOH+ are 3949.9 cm for OH stretching, 747.4 cm-1 for CaO stretching, and 481.2 cm-1 for doubly degenerate bending. The frequency for the OH mode is slightly larger as compared to our theoretical value of 3748 cm-1 and the experimental value of 3737.8 cm-1 Huber and Herzberg (1979) for a free OH. Additionally, using the finite-field approach, we predict the equilibrium permanent electric dipole moment of CaOH+ with respect to its center of mass to be 6.3 debye. This relatively large dipole moment originates from the partially charge-transferred nature of the ground-state CaOH+, which can be formally described as Ca2+OH-.
Under room temperature (300K), the molecular ion is likely to be in its vibrational ground state (97% for CaO stretching). Using the CCSD(T) method, we predict that the dissociation energy of CaOH+ in the ground rovibrational level () into and OH () is 37807 cm-1, which corresponds to 264.5 nm. This value originates from 38763 cm-1 of the electronic contribution corrected by cm-1 of the zero point energy (ZPE) difference within the harmonic approximation, where the ZPE for CaOH+ and OH is 2830 cm-1 and 1874 cm-1, respectively.
Figure 7 presents the ground and excited electronic states of the CaOH+ in the linear geometry as a function of the distance between Ca and the center of mass of OH. OH is described within the rigid-rotor approximation with the experimental internuclear distance of = 1.8332 bohr Drouin (2013). The one-photon dissociation of CaOH+ is driven by the allowed electric dipole transition from the ground state to the excited state, which is slightly repulsive around the ground-state equilibrium geometry and dissociates into Ca+ and OH. The corresponding vertical transition moment obtained with the MRCISD method is 0.8 debye. The photodissociation cross sections are given by the convolution of the transition moment function and the overlap of the ground rovibrational and excited continuum nuclear wave functions. Exact values of cross sections are sensitive to the precise location of the classical turning point of the state, which is, however, robustly predicted by the employed ab initio methods to be within the width of the ground rovibrational wave function.
VII Discussion
@@
In summary, we measured the photodissociation spectrum of single trapped CaOH+ molecular ions sympathetically cooled in Ca+ ion crystals at nm and nm. The photodissociation processes at wavelengths in these two ranges are verified to originate predominantly from a single-photon and a two-photon dissociation channel, respectively. While the one-photon dissociation cross section started to drop off around nm, a similar decrease was observed in the two-photon dissociation spectrum around nm. This implies that both processes involve the excitation of the molecular ion to its unbound first electronic excited state which leads to its dissociation. Compared to the previous workOkada et al. (2006) on single-photon dissociation of CaOH+ by Okada et al. in 2006, we report a more precise measurement of the dissociation cross sections at different wavelengths which is consistent with the previous estimation of at nm. The measured single-photon dissociation spectrum is also consistent with the photodissociation of CaOH+ at 272 nm observed by Hansen et al.Hansen et al. (2012).
@@
As for the two-photon dissociation of CaOH+ around nm, the dissociation process is found to be significantly less efficient compared to the single-photon dissociation process, as expected since higher peak intensity is typically required for higher-order processes. Yet femtosecond lasers at such wavelengths have proven to be alternatives to continuous wave lasers with nm for photodissociation of the molecular ion. The question arises as to whether nanosecond lasers that are used for ablation loading in trapped atomic ion experiments at nmPogorelov et al. (2021) are able to dissociate CaOH+. Achieving a mean photodissociation time of 1 minute for a single CaOH+ ion is estimated to require an average power of W, considering a laser with kHz, ns, and the light focused onto the ions with a beam waist of µm.
We employed advanced ab initio molecular electronic structure methods to characterize theoretically the ground-state properties of CaOH+ and to investigate the spectrum of excited electronic states involved in the measured photodissociation. We predicted the photodissociation threshold for CaOH+Ca++OH at 265 nm in very good agreement with the one-photon experimental observations. Efficient photodissociation is possible because of the coincidence of the slightly repulsive part of the excited electronic state close to its inner classical turning point, which is accessible by a decent electric dipole transition moment.
@@
The reported measurement is implemented on single CaOH+ molecular ions in small trapped ion crystals, which enables the relatively precise evaluation of the photodissociation cross section due to the repeatable positioning of the molecule. The scale of the system is also crucial for investigating a two-photon dissociation process that requires tightly focused light to achieve sufficient intensity. The reported photodissociation spectrum can be a basis for studying the rovibrational structure of CaOH+ via dissociation-based spectroscopy. Designed to investigate the interaction between femtosecond laser pulses and a linear ion crystal with a single CaOH+, this experiment represents a precursor for performing quantum logic spectroscopy with polyatomic molecular ions Schindler (2019).